A 160 gg block hangs from a spring with spring constant 7.0 N/m . At t=0s the block is 20 cm below the equilibrium point and moving upward with a speed of 125 cm/s .

What is the block's oscillation frequency?

What is the block's distance from equilibrium when the speed is 39 cm/s ?

What is the block's distance from equilibrium at t=6.0s

Show work please.
Answer #1

m= 1609 K-7Nm at t=0; xo= 20cm ; Vo = 125 cmo using, a= A Cos (t+S); Putto 20 cm = A COOS - Now ,v= - A ce Suillettd) But toе) к = m G 7 = 0.16 + 27 8² f=11.053 Hz coscillation frequency) V – A 6 si (6t+8) 39 ans =- 27.517 6.6143 Sun (at - 43.38) -